A 1.9 kg mass and a 4.4 kg mass are connected by a massless string over a pulley that is in the shape of a solid wheel having a radius 0.4 m and mass 4.7 kg. Determine the linear acceleration of the blocks. Answer in "m/s^2"
Answer #1

val 12V Page ma) my Imag Writing equation for me and m2 using the Newton I lova_F=ma T - Mg g = mia i Mag-T2 = mza — ii) addialso Pere Torque is given as, where, I is moment = I. a of mentia, and a ia angular (T2-T₂) R = It for disc, I = 1 MR² [momenа mitma + M g (ma-m,) = ma-mo) g (m, + m₂ + M) where, m, ; Ma are mass of objects g 4 acceleration due to grauiga M mass of P

Having any problem please comment.

Thank you.